OpenStudy (anonymous):

The half-life of a certain radioactive material is 71 hours. An initial amount of the material has a mass of 722 kg. Write an exponential function that models the decay of this material. Find how much radioactive material remains after 17 hours. Round your answer to the nearest thousandth. a. 0.103 kg b. 611.589 kg c. 0.414 kg d. 0 kg

OpenStudy (e.cociuba):

Use the radioactive decay formula: A = Ao*2^(-t/h): where A = resulting amt after t time Ao = initial amt (t=0) t = time of decay h = half-life of substance When t = 17 hrs A = 722*2^(-17/71) find 2^(-17/71) with a calc A = 722*.847076 A = 611.59 kg after 17 hrs

OpenStudy (e.cociuba):

So your answers going to be B :)

OpenStudy (anonymous):

thank you!

OpenStudy (e.cociuba):

No problemo! :)